# TEASMA: A Practical Approach for Test Assessment of Deep Neural Networks using Mutation Analysis

### v2

@@ VIII Conclusion
In this paper, we propose TEASMA, a practical approach for assessing the adequacy of test sets for DNNs. TEASMA is based on mutation analysis during the training phase. Such assessments are important in practice, regardless of the test set selection strategy, in order to determine whether test results can be trusted to validate DNN models.
Our results show that, although there is a strong relationship between Mutation Score (MS) and Fault Detection Rate (FDR), the shape of the relationship is typically not linear and greatly varies across DNN models. Consequently, MS is not an accurate surrogate measure for FDR and cannot be used directly as a test adequacy criterion. TEASMA’s strategy is therefore to provide a means to accurately predict the FDR of a test set from its MS so as to confidently decide whether it is adequate to provide reliable test results. If deemed adequate, the test set is then labeled and executed for validating the DNN model. Such a procedure also prevents the potential waste of effort associated with labeling test inputs that may turn out to be inadequate.
TEASMA relies on post-training mutation operators as they require much less computational resources compared to pre-training operators. Post-training operators do not require retraining and are therefore practically applicable for test adequacy assessment in all situations. Though not representative of real faults, our results show that such operators can be used for such assessments since we can accurately predict FDR based on MS.
TEASMA relies on regression analysis to build an accurate FDR prediction model based on MS using the training set. Based on training sets, prediction models yielded a very high $R^{2}$ ($R^{2}\geq 0.98$) across subjects and thus enabled TEASMA to confidently predict the FDR of an input set from its MS. Our results further indicate that such prediction models can accurately predict a test set’s FDR— with $R^{2}\geq 0.94$ and a slope close to 1—which can then reliably be used to decide about the adequacy of a test set.
According to our findings, TEASMA is thus a reliable instrument to ensure the proper validation of a DNN before deployment, especially in more critical contexts where trustworthiness is crucial.
@@
In our future work, we intend to rely on TEASMA to automate the generation of adequate test sets for DNN models.
@@ Acknowledgements
This work was supported by a research grant from Huawei Technologies Canada Co., Ltd, as well as by the Mitacs Accelerate Program. Additional support was provided by the NSERC Discovery and Canada Research Chair programs. The experiments conducted in this work were enabled in part by support provided by the Digital Research Alliance of Canada  [^1] https://alliancecan.ca.

### v3

@@ VII Conclusion
In this paper, we propose TEASMA, a practical methodology for assessing the adequacy of unlabeled test sets for DNNs. Such assessments are important in practice, regardless of the test set selection strategy, in order to determine whether high-accuracy test results can be trusted to validate DNN models. Once a test set is deemed adequate, it can then be labeled and executed.
Our empirical analysis investigated the relationship between the Adequacy Scores (AS) of test sets, calculated using four SOTA test adequacy metrics, including Distance-based Surprise Coverage (DSC), Likelihood-based Surprise Coverage (LSC), Input Distribution Coverage (IDC), and Mutation Score (MS) calculated based on post-training mutation operators. Results show a strong correlation between AS and the Fault Detection Rate (FDR) for all these metrics. However, the shape of the relationships is typically not linear and greatly varies across DNN models. Consequently, these metrics are not accurate surrogate measures for FDR and cannot be used directly for test adequacy assessment. TEASMA’s methodology provides a procedure to accurately predict the FDR of an unlabeled test set from its AS, relying on a selected adequacy metric, so as to confidently decide whether a test set can provide reliable test results. TEASMA is based on regression analysis using the training set and can be deployed with any test adequacy assessment metric for DNNs.
We conducted an extensive empirical evaluation using multiple DNN models and input sets, including large input sets such as ImageNet, with four SOTA adequacy metrics: MS, DSC, LSC, and IDC. Our results indicate that, unlike LSC, prediction models built using MS, DSC, and IDC can accurately predict a test set’s FDR— with $R^{2}$ values of 0.94 for MS and 0.90 for DSC and IDC, and a slope close to 1—which can then reliably be used to decide about the adequacy of a test set. Additionally, the average Root Mean Square Error (RMSE) between predicted and actual FDR values when using MS is 0.09, indicating a higher prediction accuracy than that of DSC and IDC, with RMSE values of 0.17 and 0.18, respectively. Therefore, MS based on post-training operators shows to be the most accurate adequacy metric when deployed within TEASMA. Overall, TEASMA is thus a reliable methodology to ensure the proper validation of a DNN before deployment, a particularly important task in critical contexts.
@@
Given the practicality of TEASMA and its ability to accurately predict FDR, we intend to rely on it, in future work, to guide the automated generation of adequate test sets for DNN models.
@@ Acknowledgements
This work was supported by a research grant from Huawei Technologies Canada Co., Ltd, as well as by the Mitacs Accelerate Program. Additional support was provided by the NSERC Discovery and Canada Research Chair programs, and the Science Foundation Ireland grant 13/RC/2094-2. The experiments conducted in this work were enabled in part by support provided by the Digital Research Alliance of Canada  [^1] https://alliancecan.ca.